Assume that z-scores are normally distributed with a mean of 0 and a standard deviation of 1. If P ( − 0.6 < z < b ) = 0.3456 , find b. b = (Round to three decimal places.) Do not round until the very last step. Any intermediate steps should include all decimal places.
Answer #1

P(-0.6<Z<b)=0.3456

=> P(Z<b)-P(Z<-0.6)=0.3456

=> P(Z<b)-0.2742531=0.3456

=> P(Z<b)=0.3456+0.2742531=0.6198531=P(Z<0.305095)

So,b=0.305095